# Routing and wavelength assignment in quantum key distribution networks: power control heuristics for quantum-classical multiplexing

### v1

@@ IV System Model
We can give a formal description of the QKD networks we want to study under a few reasonable assumptions. First we will describe the exact problem we want to solve and then we will choose a relevant scenario where we perform our numerical experiments.
IV-A Problem Statement
@@
We consider an arbitrary physical topology of an optical network as a directed graph $G(N,L,W_{T},W_{Q})$ composed of $N$ nodes and $L$ optical fiber links. $W_{T}$ is the total number of wavelengths in each link, while $W_{Q}$ is the number of reserved wavelengths for the quantum channel. We consider quantum requests $R(s,d)$ where $s$ is the source node and $d$ is the destination node. In the experiments, for each quantum request we require that four channels are established: the quantum channel, two classical control channels (one from $s$ to $d$ and one from $d$ to $s$), and an additional classical channel for user data transmission.
@@
We want to provide a valid lightpath for all the traffic requests for the given topology. In our hybrid networks, the signals in the classical channels can disrupt the quantum transmission. We will propose RWA heuristics using power control over the classical channels to minimize these interferences from the classical into the quantum lightpaths and compare them to the usual classical heuristics.
IV-B Network Assumptions
@@
We assume a QKD software-defined network (QKD SDN) with a classical control plane which uses the proposed algorithms to evaluate and allocate network resources to all the connection requests.
For our experiments, we generate random network topologies with a minimum of $8$ nodes and a maximum of $12$ nodes. The topologies are created so that any node has at least one neighbour to avoid isolated nodes in the graph. Two nodes are physically connected with a given link probability. When two adjacent nodes are connected, we set two unidirectional links (one from $a$ to $b$ and one from $b$ to $a$). We have set relatively short links of tens of kilometers due to restrictions in the achievable secure distance for commercial QKD links.
@@
The performance of the algorithms is measured in terms of blocking probability, which is the ratio between the rejected connections and total demands. Connections for both quantum and classical requests are rejected if the algorithm cannot find a path and a wavelength between the source and destination nodes. Additionally, QKD lightpaths should be established only if they allow the generation of a secret key that meets a certain quality parameter, which we measure in terms of Quantum Signal-to-Noise Ratio (QSNR), which will be introduced in the following subsection. Furthermore, if any classical channel, either associated with the requested quantum channel or with a purely classical demand causes noise to any established quantum channel that introduces more errors than the protocol can tolerate, that connection should not be established, since it will cause a failure in the key generation procedure while wasting network resources.
@@
IV-C Transmission Power Control and RWA Algorithms in QKD Networks
IV-C1 Transmission Power Control
@@
We adapt RWA algorithms described in [10] by adding an adaptive power launch feature for the classical channels. The RWA algorithms find a candidate path and available wavelength and calculate the power launch for that particular path, so that the signal satisfies a given signal-to-noise ratio (SNR) at the receiver, following a similar procedure to [11, 12]. The SNR equation can be expressed as:
$\text{SNR}=\frac{S_{c}}{N}=\frac{10^{-\alpha_{c}L_{c}}P_{tx,c}}{N}$ (1)
where $\alpha_{c}$ is the attenuation at the transmission wavelength, $L_{c}$ is the total length of the candidate path, $P_{tx,c}$ is the launch power of the classical signal and $N$ is the noise term. All the signal and noise powers will be normalized to a reference level. From this expression, we can calculate the launch power as:
$P_{tx,\ c}=\frac{SNR\cdot N}{10^{-\alpha_{c}L_{c}}}.$ (2)
@@
With these calculations, the control plane can adjust the power of the classical signal at each source transmitter to a level that both guarantees a clear reception and minimizes interference into the quantum channels.
IV-C2 RWA Algorithms
@@
We study three different heuristics, both with and without power control.
@@
  • Minimum Quantum Distance Overlap (MQDO): This algorithm identifies all possible paths between a pair of nodes, marking the links of each path that are shared with established quantum channels. Then, the algorithm attempts to allocate an available wavelength on the path with the lower accumulated shared distance using the first-fit technique. If the allocation fails, the algorithm proceeds with the next path in ascending order of accumulated shared distance with established quantum channels and tries again. The process continues until finding a suitable wavelength or until all paths are exhausted. If the allocation is not possible, the connection is blocked.
@@
  • Minimum Quantum Classical Channel Overlap (MQCCO): Similar to MQDO, the algorithm calculates the accumulated shared distance with quantum connections for each possible path, considering other established classical channels sharing any link of the path. Specifically, for each established quantum channel, if a link of the candidate path is marked as shared, and that link is also traversed by an established classical channel, the contribution to the accumulated shared distance is doubled. After that, the algorithm proceeds as described in MQDO. It tries to allocate an available wavelength to a candidate path using the first-fit technique. If the allocation is not possible, the algorithm moves to the next path in increasing order of shared distance. If the algorithm is unable to allocate resources after evaluating all candidate paths, the request is blocked.
@@
  • Quantum Totally Disjoint (QTD): The algorithm attempts to allocate a path and an available wavelength that do not share links with any established quantum channel. If the algorithm cannot find such a path or available spectral resources, the connection is blocked.
@@
For the quantum channel, all the proposed algorithms verify if the selected path meets the requirements of QSNR. In the case of quantum requests, if the resulting QSNR falls below a certain threshold, the algorithms sequentially evaluate the next path until they identify a path with a free wavelength that satisfies the QSNR threshold or block the request otherwise.
@@
Conversely, the establishment of classical channels must not degrade the performance of existing quantum connections. Hence, the RWA algorithms ensure that the selected path does not take the QSNR of established quantum channels below the specified QSNR threshold. If a path fails to meet this criterion, the algorithms proceed to assess subsequent paths until finding one that does not degrade the established quantum channels. If no such path is found, the connection is blocked. The QSNR is estimated with:
$\text{QSNR}\ =\frac{10^{-\alpha_{q}L_{q}}P_{tx,q}}{N_{F}+\gamma_{nl}\sum_{i}P_{tx,i}\cdot e^{-\alpha_{c}L_{i}}}$ (3)
where $\alpha_{q}$ is the attenuation of the fiber at the transmission wavelength of the quantum signal, $L_{q}$ is the length of the route allocated to quantum channel $q$, $P_{tx,q}$ is the transmitted power of the quantum channel, $N_{F}$ is a fixed noise term, $\gamma_{nl}$ is a fixed nonlinear coefficient term that can be adapted to the physical parameters of the fiber used in the studied network, $P_{tx,i}$ represents the classical signal launch power at the beginning of the shared span $i$, with shared length $L_{i}$ and $\alpha_{c}$ represents the attenuation coefficient at the transmission wavelength of the classical channel. We use two metrics to calculate the shared span:
  1. 1.
    Actual shared length: $L_{i}$ represents the actual length of the shared fiber.
  2. 2.
    Effective length: if $L$ represents the actual distance in span $i$ for which the power can be assumed as constant, $L_{i}$ will be set to the effective distance which is computed as [13]:
    $L_{i}=L_{\text{eff}}=\frac{1-e^{-\alpha_{c}L}}{\alpha_{c}},$ (4)
for an $\alpha_{c}$ given in natural units. This distance takes into account that the attenuation of the classical signal during propagation implies a reduced number of noise photons created by nonlinear effects. $L_{\text{eff}}$ is the distance which produces the equivalent nonlinear effects for a fixed signal power.
V Simulation Set up and Results
V-A Set up
We investigate the effects of the transmission power protocols on $10$ randomly generated topologies. The number of nodes for each topology is randomly chosen between $5$ and $10$. Each pair of nodes is connected via two unidirectional fibers with a probability of $0.5$. To ensure network connectivity, each node must have at least two neighbours. The length of the fiber links are randomly chosen between $10$ and $20$ km. Examples of the generated topologies are shown in figures 1,2.
Figure 1: Example of a random topology with $6$ nodes generated in this study.
Figure 2: Example of a random topology with $9$ nodes generated in this study.
[…]

### v2

@@ 4 Network Model
We can give a formal description of the QKD networks under study with a few reasonable assumptions. First, we will describe the exact problem we want to solve and then we will choose a relevant scenario to perform our numerical experiments.
4.1 Problem Statement
@@
We consider an arbitrary physical topology of an optical network as a directed graph $G(N,L,W_{T},W_{Q})$ composed of $N$ nodes and $L$ optical fiber links. $W_{T}$ is the total number of wavelengths in each link, while $W_{Q}$ is the number of reserved wavelengths for the quantum channel. We consider quantum requests $R(s,d)$ where $s$ is the source node and $d$ is the destination node. In the experiments, four channels are required to be established for each quantum request: the quantum channel, two classical control channels (one from $s$ to $d$ and one from $d$ to $s$), and an additional classical channel for user data transmission.
@@
We want to provide a valid lightpath for all the traffic requests for the given topology. In our hybrid networks, the signals in the classical channels can disrupt the quantum transmission. We will propose RWA heuristics using power control over the classical channels to minimize the photons the classical signals generate into the quantum lightpaths and compare them to the usual classical heuristics.
4.2 Network Assumptions
@@
We assume a QKD software-defined network (QKD SDN) with a classical control plane which uses the proposed algorithms to evaluate and allocate network resources for each connection request.
For our experiments, we generate random geometric graphs corresponding to spatial networks on an Euclidean plane. We generate network topologies with differente numbers of nodes using two different graph models: Gabriel [40] and Waxman [39], both of which capturing some of the properties of typical communication networks. Each generated graph is constructed such that every edge becomes a bidirectional connection between two nodes, represented by two unidirectional links (i.e. one link from node $a$ to node $b$ and another from $b$ to $a$). Due to restrictions in the achievable secure distance for commercial QKD links, we have set relatively short links of tens of kilometers, ensuring that the longest path in the network does not exceed 60 km. To efficiently estimate the network diameter under these constraints—especially when computing exact longest paths is computationally expensive—we approximate it by selecting the longest path among the $k$-shortest paths between all node pairs.
@@
The performance of the algorithms is measured in terms of blocking probability, which is the ratio between the rejected connections and total demands. Connections for both quantum and classical requests are rejected if the algorithm cannot find a path and a wavelength between the source and destination nodes. Additionally, QKD lightpaths should be established only if they allow the generation of a secret key that meets a certain quality parameter, which we measure in terms of the Quantum Signal-to-Noise Ratio (QSNR). Furthermore, if any classical channel, either associated with the requested quantum channel or with a purely classical demand, causes noise to any established quantum channel and that noise introduces more errors than the protocol can tolerate, that connection should not be established, since it will cause a failure in the key generation procedure while wasting network resources.
@@
5 Transmission Power Control and RWA Algorithms in QKD Networks
5.1 Transmission Power Control
@@
We adapt the RWA algorithms described in [30] by adding an adaptive power launch feature for the classical channels. The RWA algorithms find a candidate path and available wavelength and calculate the power launch for that particular path, so that the signal satisfies a given signal-to-noise ratio (SNR) at the receiver, following a similar procedure to [37, 38]. Once we have a route and know the fiber length $L_{c}$ the classical signal must travel, from the SNR expression in Equation (2), we can calculate the launch power so that the SNR at the receiver matches the quality threshold in our network as
$P_{tx,\ c}=\frac{SNR\cdot N}{e^{-\alpha_{c}L_{c}}}.$ (8)
@@
The control plane can use these calculations to adjust the power of the classical signal at each source transmitter to a level that both guarantees a clear reception and minimizes interference into the quantum channels (see Algorithm 1).
Algorithm 1 Compute Launch Power to Meet SNR Threshold
1:global classicalAttclassicalAtt
2:procedure ComputeLaunchPower(snrThresholdsnrThreshold, ClassicalPathClassicalPath)
3:  $totalAttenuation\leftarrow 0$
4:  $length\leftarrow 0$
5:  for $link\in ClassicalPath$ do
6:    $length\leftarrow length+link.length$
7:  $totalAttenuation\leftarrow e^{-classicalAtt\cdot length}$
8:  $launchPower\leftarrow snrThreshold/totalAttenuation$
5.2 RWA Algorithms
@@
We study three different heuristics for routing the classical signals, both with and without power control. In the power control version, when we set a path, we set the power of the transmitter following Equation (8).
  • 1.
@@
    Minimum Quantum Distance Overlap (MQDO): This algorithm identifies all possible paths between a pair of nodes, marking the links of each path that are shared with established quantum channels. Then, the algorithm attempts to allocate an available wavelength on the path with the lower accumulated shared distance using the first-fit technique. If the allocation fails, the algorithm proceeds with the next path in ascending order of accumulated shared distance with established quantum channels and tries again. The process continues until finding a suitable wavelength or until all paths are exhausted. If the allocation is not possible, the connection is blocked. The pseudocode for this method is shown in Algorithm 2, which uses Algorithm 3 as a subroutine. We also present here the procedure used in all the heuristics to check whether a new classical channel will reduce the QSNR of any existing quantum channel below the quality threshold (Algorithm 4) and the functions used in the calculation. Algorithm 5 computes the QSNR of a quantum channel after we add a new classical channel using Equation (3), with Algorithm 6 providing the Raman noise from the classical channels in the denominator. In the heuristics, whenever we have a request for a new classical connection, we will check if the QSNR of any existing quantum channel goes below a global quality threshold. If this is the case, that path is not selected. The connection request is blocked if there are no valid paths to establish the new classical channel.
    Algorithm 2 Minimum Quantum Distance Overlap (MQDO)
    1:procedure Minimum Quantum Distance Overlap(sourcesource, destinationdestination, numWavelengthsnumWavelengths, candidatePathListcandidatePathList, channelStatuschannelStatus, allocatedQuantumPathsallocatedQuantumPaths, edgeLengthMatrixedgeLengthMatrix)
    2:  $candidatePathListCopy\leftarrow$ copy(candidatePathListcandidatePathList)
    3:  sort(candidatePathListCopycandidatePathListCopy, Distance Overlap)
    4:  for pathpath in candidatePathListCopycandidatePathListCopy do
    5:    if meetsQSNR(pathpath) then
    6:     $w\leftarrow 0$
    7:     while $(w<numWavelengths)$ do
    8:      $available\leftarrow\textbf{True}$
    9:      for $link\in path$ do
    10:         if $channelStatus[link][w]=1$ then
    11:         $available\leftarrow\textbf{False}$
    12:      if availableavailable then
    13:         return $path,w$
    14:      else
    15:         $w\leftarrow w+1$
    16:  return [],1[],-1
    Algorithm 3 Distance Overlap (used in MQDO)
    1:global edgeLengthedgeLength
    2:global graphNetworkEdgesgraphNetworkEdges
    3:global allocatedQuantumPathsallocatedQuantumPaths
    4:procedure Distance Overlap(pathpath)
    5:  $length\leftarrow 0$
    6:  for $link\in path$ do
    7:    for keykey in allocatedQuantumPathsallocatedQuantumPaths do
    8:     for $qPath$ in $allocatedQuantumPaths[key]$ do
    9:      if $link\in qPath$ then
    10:         $length\leftarrow length+edgeLength[link]$
    11:  return lengthlength
    Algorithm 4 Meet QSNR
    1:global quantumChannelsquantumChannels
    2:global qsnrThresqsnrThres
    3:procedure meetQSNR(classicalPathclassicalPath)
    4:  for $quantumPath\in quantumChannels$ do:
    5:    if $\textsc{getQsnr}(quantumPath,classicalPath)<qsnrThres$ then
    6:     return FalseFalse
    7:  return TrueTrue
    Algorithm 5 QSNR calculation
    1:global graphNetworkEdgesgraphNetworkEdges
    2:global edgeLengthedgeLength
    3:global quantumAttquantumAtt
    4:global etaeta
    5:global fixedNoisefixedNoise
    6:function getQSNR($quantumPath,NewClassicalPath$)
    7:  $length\leftarrow 0$
    8:  for all $link\in quantumPath$ do
    9:    $s\leftarrow graphNetworkEdges[link][0]$
    10:    $d\leftarrow graphNetworkEdges[link][1]$
    11:    $length\leftarrow length+edgeLength[s][d]$
    12:  $transmitivity\leftarrow\exp(-quantumAtt\cdot length)$
    13:  $launchPower\leftarrow eta\cdot transmitivity$
    14:  $sharedNoise\leftarrow\textsc{CalculateNoise}(quantumPath,NewClassicalPath)$
    15:  $result\leftarrow\frac{transmitivity\cdot launchPower}{fixedNoise+sharedNoise}$
    16:  return resultresult
    Algorithm 6 Calculate Noise due to Classical Channels
    1:global quantumChannelsquantumChannels
    2:global classicalChannelsclassicalChannels
    3:global graphNetworkEdgesgraphNetworkEdges
    4:global edgeLengthedgeLength
    5:global classicalAttclassicalAtt
    6:global quantumAttquantumAtt
    7:global nonlinearParamnonlinearParam
    8:function CalculateNoise($quantumPath,newClassicalPath$)
    9:  $RamanNoise\leftarrow 0$
    10:  for all $classicalPath\in classicalChannels$ do $\triangleright$ Existing channels
    11:   for $i\leftarrow 0$ to length(classicalPath)1length(classicalPath)-1 do
    12:     $link\leftarrow classicalPath[i]$
    13:     if $link\in quantumPath$ then
    14:      $(s,d)\leftarrow graphNetworkEdges[link]$
    15:      $linkLength\leftarrow edgeLength[s][d]$
    16:      $power\leftarrow classicalPath.LinkPower[i]$
    17:      $RamanNoise\leftarrow RamanNoise+power\cdot\left(e^{-classicalAtt\cdot linkLength}-e^{-quantumAtt\cdot linkLength}\right)$
    18:  for $i\leftarrow 0$ to length(newClassicalPath)1length(newClassicalPath)-1 do $\triangleright$ New channels
    19:   $link\leftarrow newClassicalPath[i]$
    20:   if $link\in quantumPath$ then
    21:     $(s,d)\leftarrow graphNetworkEdges[link]$
    22:     $linkLength\leftarrow edgeLength[s][d]$
    23:     $power\leftarrow newClassicalPath.LinkPower[i]$
    24:     $RamanNoise\leftarrow RamanNoise+power\cdot\left(e^{-classicalAtt\cdot linkLength}-e^{-quantumAtt\cdot linkLength}\right)$
    25:  return $nonlinearParam\cdot RamanNoise$
  • 2.
@@
    Minimum Quantum Classical Channel Overlap (MQCCO): Similar to MQDO, the algorithm calculates the accumulated shared distance with quantum connections for each possible path, considering other established classical channels sharing any link of the path. Specifically, for each established quantum channel, if a link of the candidate path is marked as shared, and that link is also traversed by an established classical channel, the contribution to the accumulated shared distance is doubled. After that, the algorithm proceeds as described in MQDO. It tries to allocate an available wavelength to a candidate path using the first-fit technique. If the allocation is not possible, the algorithm moves to the next path in increasing order of shared distance. If the algorithm is unable to allocate resources after evaluating all candidate paths, the request is blocked. Algorithm 7 shows the pseudocode for MQCCO, which uses Algorithm 8 as a subroutine.
    Algorithm 7 Minimum Quantum-Classical Channel Overlap (MQCCO)
    1:procedure Minimum Quantum-Classical Channel Overlap(sourcesource, destinationdestination, numWavelengthsnumWavelengths, candidatePathListcandidatePathList, channelStatuschannelStatus, allocatedQuantumPathsallocatedQuantumPaths)
    2:  $sortedCandidatePathList\leftarrow$ copy(candidatePathListcandidatePathList)
    3:  sort(sortedCandidatePathListsortedCandidatePathList, Channel Overlap)
    4:  for pathpath in sortedCandidatePathListsortedCandidatePathList do
    5:    if meetsQSNR(pathpath) then
    6:     $w\leftarrow 0$
    7:     while $(w<numWavelengths)$ do
    8:      $available\leftarrow\textbf{True}$
    9:      for $link\in path$ do
    10:         if $channelStatus[link][w]=1$ then
    11:         $available\leftarrow\textbf{False}$
    12:      if availableavailable then
    13:         return $path,w$
    14:      else
    15:         $w\leftarrow w+1$
    16:  return [],1[],-1
    Algorithm 8 Channel Overlap (used in MQCCO)
    1:global edgeLengthedgeLength
    2:global graphNetworkEdgesgraphNetworkEdges
    3:global allocatedQuantumPathsallocatedQuantumPaths
    4:global channelStatuschannelStatus
    5:procedure Channel Overlap(pathpath)
    6:  $length\leftarrow 0$
    7:  for linklink in pathpath do
    8:    $occupiedWavelengths\leftarrow$ sum($channelStatus[link]$)
    9:    for keykey in allocatedQuantumPathsallocatedQuantumPaths do
    10:     for $qPath$ in $allocatedQuantumPaths[key]$ do
    11:      if $link\in qPath$ then
    12:         $length\leftarrow length+occupiedWavelengths\times edgeLength[link]$
    13:  return lengthlength
  • 3.
@@
    Quantum Totally Disjoint (QTD): The algorithm attempts to allocate a path and an available wavelength that do not share links with any established quantum channel. If the algorithm cannot find such a path or available spectral resources, the connection is blocked. The pseudocode is shown in Algorithm 9.
Algorithm 9 Quantum Totally Disjoint (QTD)
1:procedure Quantum Totally Disjoint(sourcesource, destinationdestination, numWavelengthsnumWavelengths, candidatePathListcandidatePathList, channelStatuschannelStatus, allocatedQuantumPathsallocatedQuantumPaths)
2:  for pathpath in candidatePathListcandidatePathList do
3:    if meetsQSNR(pathpath) then
4:     $search\leftarrow\textbf{True}$
5:     $foundPath\leftarrow\textbf{True}$
6:     while searchsearch do
7:      for keykey in $allocatedQuantumPaths.keys()$ do
8:         for $qPath$ in quantum_paths[key]quantum\_paths[key] do
9:         for $link\in path$ do
10:           if $link\in qPath$ then
11:            $search\leftarrow\textbf{False}$
12:            $foundPath\leftarrow\textbf{False}$
13:            break
14:      $search\leftarrow\textbf{False}$
15:    if foundPathfoundPath then
16:     $w\leftarrow 0$
17:     while $(w<numWavelengths)$ do
18:      $available\leftarrow\textbf{True}$
19:      for $link\in path$ do
20:         if $channelStatus[link][w]=1$ then
21:         $available\leftarrow\textbf{False}$
22:      if availableavailable then
23:         return $path,w$
24:      else
25:         $w\leftarrow w+1$
26:  return [],1[],-1
@@ 5.3 Quantum channel evaluation
Quantum channels are assigned using a $k$-shortest-path-first-fit heuristic. Losses are higher in the O band and we prefer the paths with shorter lengths (higher transmission). For the quantum channel, all the proposed algorithms verify if the selected path meets the requirements of QSNR. In the case of quantum requests, if the resulting QSNR falls below a certain threshold, the algorithms sequentially evaluate the next path until they identify a path with a free wavelength that satisfies the QSNR threshold or block the request otherwise.
@@
Conversely, the establishment of classical channels must not degrade the performance of existing quantum connections. Hence, the RWA algorithms ensure that the selected path does not take the QSNR of established quantum channels below the specified QSNR threshold. If a path fails to meet this criterion, the algorithms proceed to assess subsequent paths until finding one that does not degrade the established quantum channels. If no such path is found, the connection is blocked. We estimate the value of the QSNR with Algorithm 5, which follows Equation (3).
Finally, for the QTD algorithms we enforce separate classical and quantum paths when we assign the quantum channels. Any path sharing a single link with a classical signal leads to a blocked request.
6 Simulation Set up and Results
6.1 Set up
We investigate the effects of the transmission power protocols on $100$ randomly generated topologies for each graph generation model. Each topology is generated with either $10$ or $20$ nodes where each node is connected to at least one neighbour. Waxman graphs are generated with parameters $\alpha=100$ and $\beta=0.45$, adding links, if necessary, to make sure the graph is connected (there are no orphan nodes or disconnected subgraphs). Examples of the generated topologies are illustrated in Figure 1.
Figure 1: Example of four random topologies generated in our simulations using Gabriel (top) and Waxman (bottom) graphs with $10$ (left) and 2020 nodes (right).
We assume $50$ wavelengths per fiber, $10$ for the quantum channels, in the O-band, which generally implements Coarse WDM with a wider channel spacing (around 20 nm) [42], and $40$ for the classical channels in the C-band, assuming Dense WDM with a 100 GHz channel separation [41]. The noise terms $N_{f}$ and $\gamma_{nl}$ are set to $10^{-4}$ and $2.25\cdot 10^{-4}$, respectively, considering dark counts dominate the fixed noise and the Raman noise model described in Section 3. We consider a normalized $P_{tx,q}$ where 1 is the power of a single photon per pulse and two fiber attenuation coefficients: $\alpha_{q}=$0.32\text{\,}\mathrm{d}\mathrm{B}\mathrm{/}\mathrm{k}\mathrm{m}$\$ around $1310\text{\,}\mathrm{nm}$ and $\alpha_{c}=$0.17\text{\,}\mathrm{d}\mathrm{B}\mathrm{/}\mathrm{k}\mathrm{m}$\$ around $1550\text{\,}\mathrm{nm}$, each corresponding to standard silica optical fiber at the usual wavelengths in the O and C bands assigned to the quantum and classical channels, respectively, in many QKD hybrid networks. Channels can be established only if a candidate path with available resources is found and the QSNR of the quantum lightpath at reception is, at least, $5\text{\,}\mathrm{dB}$.